2 Angles are Supplementary. 3 times the measure of the smaller angle is 5 less than twice the measure of the larger angle. Find the measure of each angle.
Answer by mananth(16946)

let the smaller angle be x deg
the larger angle will be 180-x
..
3x= 2(180-x)-5
3x=360-2x-5
3x+2x=355
5x=355
x=71
The other angle will be 180-x
180-71= 109 deg